### v1

# Facing Asymmetry - Uncovering the Causal Link between Facial Symmetry and Expression Classifiers using Synthetic Interventions

Abstract
Understanding expressions is vital for deciphering human behavior, and nowadays, end-to-end trained black box models achieve high performance. Due to the black-box nature of these models, it is unclear how they behave when applied out-of-distribution. Specifically, these models show decreased performance for unilateral facial palsy patients. We hypothesize that one crucial factor guiding the internal decision rules is facial symmetry. In this work, we use insights from causal reasoning to investigate the hypothesis. After deriving a structural causal model, we develop a synthetic interventional framework. This approach allows us to analyze how facial symmetry impacts a network’s output behavior while keeping other factors fixed. All 17 investigated expression classifiers significantly lower their output activations for reduced symmetry. This result is congruent with observed behavior on real-world data from healthy subjects and facial palsy patients. As such, our investigation serves as a case study for identifying causal factors that influence the behavior of black-box models.
Keywords:
Facial Expressions Facial Asymmetry Unilateral Facial Palsy Causal Inference Intervention
{NoHyper}**footnotetext: These authors contributed equally to this work.
1 Introduction
Emotional expressiveness is a crucial topic in our daily life for communicating our internal state and for understanding other people [95, 76]. The state-of-the-art for automatically classifying the six base emotions [23] is achieved by end-to-end trained black box neural networks [65, 82, 110, 102, 24, 17]. However, it remains unclear how the internal decision-making processes of these models respond to out-of-distribution inputs that were not accounted for during training. Specifically, we observe a performance degradation when classifying facial expressions in individuals with unilateral facial palsy, a condition impairing the ability to produce symmetrical facial expressions due to underlying nerve damage. Although intuition suggests that facial asymmetry could influence model behavior, we lack a quantifiable way to test and validate this hypothesis.
We leverage causal reasoning principles to address this limitation and move beyond empirical analytics to uncover a contributing factor in the underlying mechanisms driving model decision-making. Specifically, our work answers the interventional question [4]: “If we only change the facial symmetry for an input, then how does the output of an expression classifier behave?” First, we provide evidence that a symmetry bias exists for real-world data inside all models using associational methods [9, 61, 73, 71]. Second, moving up on the causal hierarchy [4], we build an interventional framework derived from a structural causal model that allows us to generate synthetic faces and connect symmetry with classifier outputs. To accurately quantify this relationship, we develop an interpretable score and an accompanying hypothesis test. As a case study, we analyze 17 expression classifiers and find significant changes in their predictions for all of them. Specifically, we find that decreases in facial symmetry result in lower logit activations. Our study highlights the importance of symmetry influencing expression classifiers, emphasizing the general need for investigations beyond predictive performance.
2 Related Work
Synthetic data has become a widely accepted tool for evaluating and training computer vision models in diverse applications, such as object detection [103, 56, 94, 97], pose estimation [94, 15, 38, 96], segmentation [81, 80, 13, 77, 92], 3D reconstruction [75, 37, 25, 69, 17], and also for facial tasks [15, 43, 7, 6, 75]. We develop a generative interventional framework that fixes possible other confounding factors to isolate the impact of facial asymmetry on expression classification.
Facial Expression Classification. Since the standardization of facial expression into six base emotions by Ekman [23], state-of-the-art performance for automated classification is achieved by end-to-end trained black-box models [1, 2, 65, 82, 12, 100, 50, 111, 17, 83, 84]. While such models reach high performance, their inner workings remain opaque. Hence, the relationship between facial symmetry and predictions remains unclear, especially in medical contexts like facial palsy [10, 8, 3, 18, 40, 42, 57, 58]. To address this uncertainty, we study the effects of facial asymmetry on expression classifiers in a controlled setting: the expression space of 3D Morphable Models [5, 21, 29, 112], more specifically FLAME [48]. While EMOCA [19], an extension of DECA, also relies on the FLAME expression space for classification, our approach takes a different route. We maximize the logit activation output for each model and emotion combination by leveraging the expression space, ensuring optimal performance. We then rely on methods from explainability to perform an in-depth investigation into the model behavior.
Explaining Model Decisions Behavior. Local explainability methods, e.g., [88, 74, 85, 87, 90], are used to investigate the behavior of machine learning models for singular inputs, e.g., highlight important image regions. Further, in [44], such local explanations are summarized to form a more conclusive general understanding of a classifier (global explanation). The focus is put on shortcut biases leading to so-called “Clever Hans predictors” [44]. However, such local attribution methods necessitate a semantic interpretation, for example, by a domain expert. Especially for medical applications, more abstract but relevant features increase the complexity [54, 78, 55]. We are interested in analyzing facial symmetry, a complex feature not directly part of the input. Global approaches, for example, [41, 73], can determine the usage of such abstract features. Especially, [73] is based on causal principles [70, 60, 64] and tests for conditional dependence between the feature and the network predictions given the labels. In [72, 62, 61, 9], this approach is applied to various application domains such as skin lesion analysis, digital agriculture, or emotion classification. However, here we go one step beyond and extend their approach by an interventional framework to generate more in-depth answers according to Pearl’s causal hierarchy [4].
Synthetic Face Generation. Generative models have long been the go-to approach for modeling human faces. Ranging from parametric 3D Morphable Models (3DMMs) [5, 59, 29, 21], Active Appearance Models [16, 51, 34, 28], or learned in an entirely data-driven manner [48, 105, 112, 104, 29, 69, 99]. The disentanglement of identity, expression, pose, and appearance is a powerful tool for bias identification [15, 43] or image manipulation [93, 19, 52, 66]. In contrast, Generative Adversarial Networks prioritize photorealism over control, embedding multiple facial properties into a single latent representation, making it challenging to have specific control over the generation [7, 6, 39, 68, 67, 14, 63]. Many approaches utilize neural networks to compute the 3DMM parameters from 2D images, either by reconstructing faces [25, 69, 104, 105, 99, 49, 33] or training in an adversarial manner [52, 66, 93, 19]. We aim to quantify the impact of facial asymmetry on the predictive behavior of expression classifiers. To achieve this, we prioritize control over photorealism in our generative pipeline using DECA [25]. Hence, we can fix other confounding factors, like appearance, lighting, and pose, at the same time. Building upon FLAME [48], we alter the geometric face model to induce subtle variations, thereby creating realistic facial asymmetry.
3 Evaluating Models by Intervening on Facial Symmetry
Studying mimicry is crucial when analyzing facial palsy, which impacts the mobility of the facial muscles. An objective evaluation of the nerve damage is commonly done via data-driven methods [10, 32, 42, 40]. In this work, we focus on one facial feature likely impacting the downstream performance of expression classifiers: facial symmetry. We start by detailing our investigation’s causal model and framing the question we are trying to answer in Pearl’s causal hierarchy [4]. Afterward, we describe the adapted 3D Morphable Model to perform interventions by changing one face side’s geometry. Lastly, we derive an interpretable score and corresponding significance test to quantify the impact of facial symmetry on a model’s prediction.
[…]
Figure 1: Expression classifier structural causal model: $Y$ is the expression influenced by the latent distribution of all facial images (hatched box), $S_{*}$ samples from this latent distribution [73]. $\mathcal{D}_{\text{train}}$ is the training data distribution. The model architecture is an exogenous variable, and weights $\theta$ are learned using an optimizer, i.e., a training algorithm. The model’s predictions $\hat{Y}$ result from the trained model $\mathbb{F}_{\theta}$. We investigate whether $\mathbb{F}_{\theta}$ is independent of the model predictions $\hat{Y}$ ( dashed red arrow). Additionally, we analyze the changes in behavior for varying facial symmetries. Toward this goal, we perform synthetic interventions ($do(\text{Facial Symmetry}:=s)$) on facial symmetry variable using 3d morphable models $I_{\varphi^{(e)}}$. Note that these $I_{\varphi^{(e)}}$ are a part (subpopulation) of the latent distribution of all facial images. Adapted from Figure 2 in [73].
3.1 Preliminaries & Causal Point of View
Causal inference tries to answer causal questions from data [64]. This includes interventions, i.e., additional experiments, and purely observational data. Importantly, causal questions can be categorized into a hierarchy. This so-called Pearl’s causal hierarchy (PCH) [4] consists of the following three levels ordered by increasing difficulty: associational, interventional, and counterfactual questions. The latter two are analyzed using the do-operator [60], which changes a variable to a constant value, e.g., we would write $do(\text{Facial Symmetry}:=s)$ for the variable facial symmetry.
Furthermore, framing data-generating processes and complex interactions of our physical reality as directed graphs enables us to precisely define and investigate the underlying causal mechanisms [60, 64]. The resulting models are called structural causal models (SCMs), and we include a formal definition in the supplementary material. Nevertheless, to understand this framework, it is important to interpret the dependencies, i.e., connections in the graph, as assignments and not as algebraic mappings [64]. Specifically, the connections between variables in such an SCM function like physical mechanisms and not like instantaneous equations. This work extends a specific SCM to model supervised learning [73].
We visualize our SCM for expression classification in Fig. 1, enabling us to study different questions about the decision process. For example, Reimers et al. [73] answer associational questions of whether a feature, such as facial symmetry, is used during the prediction, i.e., does the red dashed arrow exist in Fig. 1. Intuitively, they measure if there is a statistically significant shift in classifier outputs for inputs of the same class but with different feature manifestations.
Other works visualize such significant changes for feature variations [61, 9]. However, they lack actionable descriptions of how the model would behave if a particular feature, e.g., facial symmetry, changes for a specific individual. In this work, we go one step up on the PCH. We employ a synthetic rendering pipeline to alter the facial symmetry while controlling other factors. Hence, we answer the interventional question: “If we change the facial symmetry for an input, then how does the output of the expression classifier behave?” Please note that the levels of the PCH are disjunct and increasing in difficulty. In [4], the authors prove the Causal Hierarchy Theorem (CHT), which states that one needs data of at least the corresponding level to answer causal questions of that level.
In the following, we describe how we generate synthetic data ($I_{\varphi^{(e)}}$ in Fig. 1), where we have fine-grained control over facial symmetry and realized emotional expressions. Using this framework, we generate new interventional data. Hence, we do not violate the CHT [4]. Further, while our approach of synthetic generation necessarily introduces a domain shift (see Fig. 1), we argue that it enables us to go beyond simple interventions. Specifically, our framework allows us to vary the facial symmetry for a specific individual and measure changes in the classifier outputs while fixing other confounding factors. Finally, we discuss how we quantify systematic output changes and determine significance.
3.2 Facial Symmetry Intervention Framework
We require a controllable face generation method to answer interventional questions of the form: “If we change the facial symmetry for an input, then how does the output of the emotion classifier behave?” Additionally, the generation process has to ensure that only facial expressions contribute to the changes measured by the expression classifier. Therefore, we select a 3D Morphable Model (3DMM) [5, 29, 59, 48, 21], to be precise FLAME [48], used in the DECA architecture to create synthetic facial images [25]. Although the generated faces introduce a domain shift, the underlying representation of identity, expression, and appearance gives us complete control over individual changes. Therefore, this disentanglement ensures we can causally link facial changes to the model’s predictive behavior. In the following, we detail our face generation framework to (a) find the expression parameters for optimal classifier activation and (b) introduce a controllable symmetry value $s$ for interventional reasoning.
For all synthetic face images $I$ in this work, we utilize the DECA pipeline [25]: $I=\mathcal{R}(\mathcal{M},\mathcal{B},c)$, composed of the face model $\mathcal{M}$, camera position $c\in\mathbb{R}^{3}$ (fixed to $[0,0,0]^{T}$ in this work) and illumination process $\mathcal{B}$ used in the differential renderer $\mathcal{R}(\cdot)$ [25]. To study facial asymmetry, we alter the face model $\mathcal{M}$ formally defined as $\mathcal{M}(\beta,\vartheta,\varphi,\alpha)=\{\mathcal{G}(\beta,\vartheta,\varphi),\mathcal{A}(\alpha)\}$. DECA employs the geometric components of FLAME $\mathcal{G}$ using the identity $\beta\in\mathbb{R}^{100}$, expression $\varphi\in\mathbb{R}^{50}$, and pose $\vartheta\in\mathbb{R}^{6}$ blendshape parameters [48]. The texture is computed from the appearance model $\mathcal{A}$ from the Basel Face Model using the parameter $\alpha$ [59, 29, 25]. In FLAME, the face geometry is modeled as
$\mathcal{G}(\beta,\vartheta,\varphi)=W(T+B_{I}(\beta,\mathcal{I})+B_{P}(\vartheta,\mathcal{P})+B_{E}(\varphi,\mathcal{E}),J(\vartheta),\vartheta,\mathcal{W}),$ (1)
with $W$ being a standard skinning function to rotate the modified $N$ face vertices of the template model $T\in\mathbb{R}^{N\times 3}$ around predefined FLAME joints $J\in\mathbb{R}^{3K}$. $B$ denotes a linear blend skinning (LBS) [45] function of the according blend shapes with identity $\mathcal{I}\in\mathbb{R}^{100\times N\times 3}$, expression $\mathcal{E}\in\mathbb{R}^{50\times N\times 3}$, and pose $\mathcal{P}\in\mathbb{R}^{6\times N\times 3}$. We use the blending weights $\mathcal{W}\in\mathbb{R}^{K\times N}$ of the original FLAME model [48].
Our changes must ensure that (a) under full facial symmetry, the original geometry holds, and (b) a symmetry scalar $s$ specifies facial symmetry and enables interventional queries. Furthermore, we formalize a time parameter $t$ to control the interpolation between neutral and a target facial expressions [49, 21].
We extend a recent approach by freezing geometry parts to simulate facial asymmetry [105]. Using a scaling parameter $s$, we can simulate different freeze states ranging from $0.0$ defining complete asymmetry to $1.0$ defining complete symmetry. We artificially induce facial asymmetry by changing only the left side of the face (person’s point of view). Therefore, we recompose the FLAME expression space such that $B_{E}(\varphi,\mathcal{E})=B_{E}(\varphi,\mathcal{E}^{L})+B_{E}(\varphi,\mathcal{E}^{R})$. Thus, we define
$\mathcal{E}^{L}_{i}=\begin{cases}\mathcal{E}_{i},&\text{if the vertex $i$ is on the left side of the face}\\ 0,&\text{otherwise}\end{cases}$ (2)
such that the linear blend skinning function $B_{E}(\varphi,\mathcal{E}^{L})$ changes only vertices on the left side of the face [45, 105]. The same applies to $\mathcal{E}^{R}$. Scaling the blendshape vectors in $\mathcal{E}^{L}$ with $s$ induces a symmetry difference between the faces’ sides. Lastly, we multiply the expression parameters $\varphi$ with $t$ to create dynamic expressions. Our geometric face model $\mathcal{G}_{s,t}$ with symmetry parameter is
$\begin{split}\mathcal{G}_{s,t}(\beta,\vartheta,\varphi,s,t)=&W(T+B_{S}(\beta,\mathcal{S})+B_{P}(\vartheta,\mathcal{P})+\\ &B_{E}(t\cdot\varphi,\mathcal{E}^{R})+B_{E}(t\cdot\varphi,s\cdot\mathcal{E}^{L}),J(\vartheta),\vartheta,\mathcal{W}).\end{split}$ (3)
(a) neutral
(b) angry
(c) disgust
(d) fear
(e) happy
(f) sad
(g) surprise
(h) $t=\nicefrac{{0}}{{6}}$
(i) $t=\nicefrac{{1}}{{6}}$
(j) $t=\nicefrac{{2}}{{6}}$
(k) $t=\nicefrac{{3}}{{6}}$
(l) $t=\nicefrac{{4}}{{6}}$
(m) $t=\nicefrac{{5}}{{6}}$
(n) $t=\nicefrac{{6}}{{6}}$
Figure 2: We display the optimized synthetic face images $I_{\varphi^{(e)}}$ for the neutral expression (2(a)) and for the six base emotions (2(b)) - (2(g)) based on the ResidualMaskingNet classifier [65]. Furthermore, we simulate with our geometric face model $\mathcal{G}_{s,t}(\cdot)$ different interpolations $t$ for a symmetry of $s=0.0$. At $t=0.0$ (2(h)) we have a neutral expression morphing into an asymmetric happy expression at $t=1.0$ (2(n)).
Thus, the synthetic face image $I_{\varphi^{(e)}}$ updates for a single individual, i.e., $\beta$ and $\alpha$ are fixed (omitted for clarity), and target expression ($e$) vector $\varphi^{(e)}$ with symmetry $s$ and temporal dynamic $t$ to $I_{\varphi^{(e)}}(s,t)=\mathcal{R}(\mathcal{G}_{s,t}(\beta,\vartheta,\varphi^{(e)},s,t),\mathcal{A}(\alpha),\mathcal{B},c)$.
Within this framework, for an individual, we can (a) modify expressions, (b) simulate facial asymmetry, and (c) simulate movements, ensuring that only changes in facial expression result in changes in the classifier’s behavior. While our synthetic faces are a domain shift for most classifiers, the comparisons are all relative and contained within this new domain. Hence, representing out-of-domain scenarios in which they are applied [84, 83, 82, 65, 7, 6]. Furthermore, we optimize the facial expression parameters such that a classifier output $\mathbb{F}^{(e)}_{\theta}$ correctly identifies the given image $I_{\varphi^{(e)}}$ as the target emotion via
$\varphi^{(e)}=\operatorname*{arg\,min}_{\hat{\varphi}^{(e)}\in[-3,3]^{|\varphi^{(e)}|}}1-\mathbb{F}_{\theta}^{(e)}(I_{\hat{\varphi}^{(e)}}).$ (4)
For this estimation problem, all parameters apart from $\varphi$ are fixed during the optimization, minimizing other confounding factors [9, 22, 101, 15], enforcing that only changes in facial expression influence the classifier output. Given that we cannot use a gradient-based optimizer as changes in $\varphi$ result in no changes in the parameters of $\mathbb{F}_{\theta}$, we use the differential evolution algorithm [89] for optimization using a search range of $[-3,3]$ [48, 21]. In Fig. 2, we visualize renderings for the six base emotions given the ResidualMaskingNet as classifier [65]. The supplementary material provides more examples and expressions parameters $\varphi^{(e)}$.
3.3 Measuring Systematic Change
Given our rendering pipeline $I_{\varphi^{(e)}}$, specified in the previous section, we need a score function to measure systematic changes in expression classifier behavior concerning facial symmetry. Hence, we define a facial symmetry impact score for a specific trained model $\mathbb{F}_{\theta}$. To be precise, we measure one score for each possible expression $e$ predicted by the selected classifier, henceforth, $\mathbb{F}_{\theta}^{(e)}$.
Using $I_{\varphi^{(e)}}$ with a sampled identity, i.e., fixed $\alpha$ and $\beta$, we generate synthetic images for timesteps $t$ and facial symmetries $s$. Now, $\mathbb{F}_{\theta}^{(e)}(I_{\varphi^{(e)}}(s,t))$ defines a surface, where for each $s$ and $t$, we have an output activation of $\mathbb{F}_{\theta}$ for emotion $e$. Figure 3(a) visualizes two of these surfaces for the neutral and happy emotion. Ideally, we would want to see no changes along the $s$ axis in these surfaces, i.e., the model is unbiased concerning symmetry. We can measure these changes by investigating the partial derivatives $\nabla_{s}\mathbb{F}_{\theta}^{(e)}(I_{\varphi^{(e)}}(s,t))$. A positive $\nabla_{s}$ indicates higher model outputs for increased symmetry, which is reversed for negative $\nabla_{s}$. An unbiased model activation surface (blue) is visualized in Figure 3(b). This optimal surface is characterized by $\nabla_{s}\mathbb{F}_{\theta}^{(e)}$ being zero for any valid $s$ and $t$.
Of course, in reality, we do not expect the outputs of any model to stay constant for changing symmetry values. Many factors can impact the model outputs, even for small visual changes. Nevertheless, we expect an unbiased model to show no systematic behavioral changes, e.g., categorically lower outputs for smaller symmetry values $s$. Hence, a more realistic ideal surface would be a noisy version of the visualization in Figure 3(b). In other words, for an unbiased model that does not change behavior for different facial symmetry, we expect that $\mathbb{E}_{s,t}[\nabla_{s}\mathbb{F}_{\theta}^{(e)}(I_{\varphi^{(e)}}(s,t))]$ is approximately zero for some joint distribution of symmetry values $s$ and timesteps $t$. Without loss of generality, let $[0,1]$ be a valid domain for $s$ and $t$ respectively, then we define our facial symmetry impact score $\mathcal{S}$ for a specific model $\mathbb{F}_{\theta}^{(e)}$ and for a fixed individual $I_{\varphi^{(e)}}$ as
[…]
where $p$ is the density function describing the joint distribution of $s$ and $t$.
Calculating $\mathcal{S}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}})$ directly is intractable. Hence, we assume that $s$ and $t$ are independent and uniformly distributed. Although this is a strong assumption, we can utilize our rendering pipeline (see Section 3.2) to ensure these conditions in our synthetic data. By doing so, we can approximate $\mathcal{S}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}})$ by evaluating $\mathbb{F}_{\theta}^{(e)}$ at a grid of finitely many equidistant samples of $I_{\varphi^{(e)}}(s,t)$.
Let $\mathfrak{T}$ and $\mathfrak{S}$ be a set of equidistant time and symmetry steps in $[0,1]$, then
$\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}})=\frac{1}{|\mathfrak{S}|\cdot|\mathfrak{T}|}\sum_{s\in\mathfrak{S}}\sum_{t\in\mathfrak{T}}\nabla_{s}\mathbb{F}_{\theta}^{(e)}(I_{\varphi^{(e)}}(s,t)),$ (6)
approximates $\mathcal{S}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}})$. To estimate the gradient on our finite grid of $\mathfrak{T}$ and $\mathfrak{S}$, we use the implementation of [26] by the library NumPy [35]. This algorithm minimizes the error between the actual gradient and the estimate at a grid position by solving a system of linear equations of the neighboring grid points.
(a) Logit activations from neutral towards happy
(b) Optimal compared to the measured activation
(c) Computed $\nabla_{s}\mathbb{F}$ of the optimal and measured surfaces
Figure 3: Visualization of our impact score for a classifier’s happy logit activation: In a synthetic setting, a model was shown a face transition from neutral to a happy expression (3(a)). A model would be invariant toward changes along the symmetry axis if $\nabla_{s}\mathbb{F}=0$. However, the actual activation logits (happy) show a lower activation (3(b)). This is more evident in the visualization of the estimated $\nabla_{s}$ in (3(c)).
While $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}})$ enables us to investigate changes for a single individual $I_{\varphi^{(e)}}$ with fixed $\alpha$ and $\beta$, we are additionally interested in explaining models $\mathbb{F}_{\theta}$ more globally concerning specific emotions. Hence, we define a global score for an emotion $e$ as $\mathcal{S}(\mathbb{F}_{\theta}^{(e)})=\mathbb{E}_{\alpha,\beta}[\mathcal{S}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}})]$. For a set of $N$ individuals $\mathfrak{I}$, by using the same assumptions as in Eq. 6, we approximate $\mathcal{S}(\mathbb{F}_{\theta}^{(e)})$ with
$\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)})=\frac{1}{N}\sum_{I_{\varphi^{(e)}}\in\mathfrak{I}}\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}}).$ (7)
3.3.1 Testing for Statistical Significance:
Values for $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)})$ and $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}})$ close to zero indicate less change for varying facial symmetry values $s$. Further, the sign of our scores can be interpreted as over- (positive) or under-predicting (negative) an emotion for increasing symmetry. However, we also need to specify at which point values of $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)})$ or $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}})$ are statistically significant.
We utilize permutation hypothesis tests [30] for this goal in which the values of $\mathbb{F}_{\theta}^{(e)}$ in our grid of synthetic inputs are shuffled. To control for the influence of $t$, i.e., the onset expression’s strength, we only shuffle values of $\mathbb{F}_{\theta}^{(e)}$ while fixing $t$. In other words, we permute along the symmetry axis in Figure 3(a). Afterward, the corresponding $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}}^{(perm.)})$ is recalculated (Eq. 6). The process repeats $K$-times to generate our distribution under the null hypothesis $H_{0}$, which is that the observed value $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}})$ is zero. By counting how often we observe permutations where $|\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}}^{(perm.)})|>|\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}})|$, we can determine a $p$-value. Note that the absolutes are needed because negative scores are valid. If the $p$-value is smaller than a significance level $\delta$, we discard $H_{0}$, i.e., the observed score $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}})$ is statistically significant. We provide the hypothesis test pseudo-code in the supplementary material (Alg. 1).
While our approach tests for significance concerning a certain $I_{\varphi^{(e)}}$ with fixed $\alpha$ and $\beta$, regarding $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)})$, we additionally apply the Holm-Bonferroni correction on our results [36]. This is a sequential correction method to control the family-wise error rate for repeated hypothesis tests. In other words, ensuring we do not overestimate significance, i.e., increase type-I errors, for a pre-specified $\delta$. Finally, we report the ratio of significant results of the corrected tests, which intuitively captures how often we observe significant changes in behavior. Note that, in contrast, $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)})$ measures how strong (and in which direction) these changes are on average. This means statistical significance is possible even if the effect size, i.e., the actual systematic change, is low. With the scores derived above combined with the corresponding statistical hypothesis tests, we can investigate the interventional question posed in Section 3.1: “If the facial symmetry for an input changes, then how does the output of the expression classifier behave?”
4 Experiments and Results
Our investigation focuses on the impact of facial symmetry on data-driven expression classifiers. Before we detail our experiments, we want to state our hypothesis clearly: We expect that most classifiers show systematic differences in their behavior when intervening on facial symmetry. Given that one face half exhibits less movement for unilateral facial palsy and our synthetic symmetry data, we assume a reduction in logit-activations for reduced symmetry based on observations in related studies [9, 106]. Because most facial expression datasets, e.g., [20, 47, 46, 53], contain mostly healthy symmetric faces. Our study is limited to the six base emotions [23], omitting neutral and contempt for comparability.
Compared to other research that requires training [43, 15], we can asses 17 existing expression classifiers. Please note that our selection of models is not intended to be comprehensive or representative of all possible classifiers but as a first case study. We focus on a subset of models that provide code and model weights, which are likely to be applied more frequently in out-of-domain scenarios like medicine and psychology. To test our hypothesis, we perform two groups of experiments. First, we investigate if a symmetry bias exists for real-world data inside classifiers using associational methods [73, 71, 61, 7]. Second, moving up on the causal hierarchy [4], we connect facial symmetry and model output using our synthetic intervention framework and score.
4.1 Experiment 1: Observations on Real-World Data
(a) Synth. neutral
(b) Synth. happy
(c) Proband neutral
(d) Proband happy
(e) Patient neutral
(f) Patient happy
Figure 4: We display two example faces per analyzed group (synthetic, probands, and patients). Both healthy probands and patients with unilateral facial palsy were instructed to mimic the shown emotions, similar to the FER2013 benchmark [20].
Although intuition suggests that facial asymmetry influences model behavior, we must first check if this bias is present in expression classifiers. We utilize associational methods [73, 71, 61, 7] to show that classifiers significantly change their behavior, i.e., the red arrow exists in Fig. 1. Specifically, we attempt to validate our hypothesis on the first level of Pearl’s hierarchy [4], using real-world data recorded on 36 healthy probands (18-67 years, 17 ♂, 19 ♀) and 36 patients (25-72 years, 8 ♂, 28 ♀) with unilateral chronic synkinetic facial palsy. Probands were recorded using a RealSense camera (Intel Corporation, Santa Clara, USA) and patients with the 3dMD face system (3dMD LLC, Georgia, USA), see Fig. 4. We model two symmetry properties: the presence of facial palsy (binary) and LPIPS [108] similarity (continuous) among face sides.
The participants’ expressions are captured while mimicking the six basic emotions four times in a random order [32, 6, 9, 7] following FER2013 [20]. We focus on the happy expression as it most impacts emotional expressiveness [76, 95, 8]. Examples of each face group are shown in Fig. 4. We measure an average classification accuracy of $97.40\%$ for probands and $58.25\%$ for patients among the FER2013 models (due to the similar recording setup), indicating the decreased performance under facial palsy. A more fine-grained analysis can be found in the supplementary material.
DAN [102] DDAMFN++ [110] HSEmotion [82] PosterV2 [50] DAN [102] DDAMFN++ [110] HSEmotion [82] PosterV2 [50] EmoNeXt-Base [24] EmoNeXt-Large [24] EmoNeXt-Small [24] EmoNeXt-Tiny [24] ResidualMaskingNet [65] Segmentation-VGG19 [98] DAN [102] DDAMFN++ [110] PosterV2 [50]
AffectNet7 AffectNet8 FER2013 RAFDB
Facial Palsy
LPIPS symmetry [108]
Table 1: Significance results ($p<0.01\rightarrow$ ) of [73] on our data for three symmetry features. We investigate happy logits with respect to the binary facial palsy state and LPIPS [108] similarity between the face sides for real-world images of facial palsy patients and healthy probands.
Following related work [9, 73, 61], we denote all significant behavior changes in Table 1 using the majority decision of three conditional independence tests [27, 79, 11] (for detailed hyperparameters see supplementary material). We find all 17 models show a statistically significant shift in their happy activations for varying facial symmetries in the real-world data. Please note that DAN [102], trained on RAFDB [47, 46], is the only classifier where we find no significance regarding binary facial palsy, which is a highly discretized form of symmetry. However, the continuous symmetry measure LPIPS [108] indicates the same behavior changes. These results provide first evidence for our hypothesis that expression classifiers are biased toward facial symmetry, especially concerning downstream applications.
Furthermore, we provide more qualitative investigations of the output changes in the supplementary material. For most classifiers, we observe, on average, a decrease in activations. Given the performance decrease, this observation is expected and indicates uncertainty in predicting the happy class. However, these are associational investigations [4], i.e., we do not isolate changes in facial symmetry from confounding factors. Hence, while we observe changes in classifier behavior on real-world data, our following interventional investigation in Section 4.2 provides more reliable, actionable insights.
4.2 Experiment 2: Synthetic Facial Symmetry Interventions
(a) FER2013 [20]
ResidualMaskingNet [65]
(b) AffectNet [53]
HSEmotion-7 [82, 83]
(c) RAFDB [47, 46]
DAN [102]
Figure 5: We display a model’s activation (mean and std.) curve at $t=1.0$ for each expression recognition dataset. Note that the x-axis is inverted, so we start with high symmetry. Lower symmetry generally results in lower logit activations across all expressions, with hatched lines indicating misclassification. We show the surface variants, such as Figure 3(a), in the supplementary material.
To confirm our hypothesis on how facial symmetry affects expression classification beyond the associational level, we perform synthetic interventions using the framework described in Sec. 3. These enable us to measure the impact of facial asymmetry and model output. Therefore, we create a population $\mathfrak{I}$ of $200$ identities sampled from a standard normal distribution (different $\alpha$ and $\beta$). Following Eq. 4, we optimize the facial expression $\varphi^{(e)}$ for each model and identity at $t=1.0$ and $s=1.0$. To apply interventions, our finite grid spans ten equidistant symmetry steps ($s\in[0,1]$) and 90-time steps ($t\in[0,1]$), simulating three-second expression onset. We then derive the mean logit surfaces (compare Fig. 3) of all classifiers over $\mathfrak{I}$ and display selected classifiers at $t=1.0$ in Fig. 5.
First, we observe that facial symmetry impacts each expression’s logit activation for the models irrespective of the dataset. The ResidualMaskinNet [65] trained on FER2013 [20] does not reach high activations for fear and sad. Further, they decrease even more for lower symmetry values. Expressions such as angry, disgust, or surprise have higher activations and seem to be affected only by more pronounced asymmetry. Especially fear also seems problematic for the other classifiers highlighted in Fig. 5. Notably, the same behavior holds for HSEmotion [82] and DAN [102]: Lower symmetry leads to lower output activations for all expressions, providing further evidence to confirm our hypothesis.
Table 2: We compute $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)})$, defined in Eq. 7, among our population $\mathfrak{I}$. The models are grouped by the training dataset, and the $\dagger$ annotates models trained by us (see the supplementary material for details); otherwise, the provided models’ weights were used respectively. All $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)})$ are significant for the majority of individuals.
Dataset $\mathbb{F}_{\theta}$ Model Angry Disgust Fear Happy Sad Surprise
AffectNet7 DAN [102] 0.0249 0.0234 0.0098 0.0242 0.0192 0.0279
DDAMFN++ [110] 0.0050 0.0035 0.0019 0.0040 0.0027 0.0037
HSEmotion [82] 0.0229 0.0308 0.0107 0.0239 0.0211 0.0268
PosterV2 [50] 0.0214 0.0192 0.0133 0.0234 0.0168 0.0290
AffectNet8 DAN [102] 0.0209 0.0222 0.0091 0.0112 0.0207 0.0281
DDAMFN++ [110] 0.0030 0.0032 0.0026 0.0028 0.0034 0.0036
HSEmotion [82] 0.0136 0.0156 0.0076 0.0111 0.0150 0.0250
PosterV2 [50] 0.0205 0.0228 0.0129 0.0214 0.0171 0.0269
FER2013 EmoNeXt-Small [24] 0.0157 0.0017 0.0078 0.0195 0.0074 0.0238
EmoNeXt-Tiny [24] 0.0092 0.0017 0.0050 0.0124 0.0051 0.0219
EmoNeXt-Base [24] 0.0089 0.0009 0.0096 0.0184 0.0096 0.0227
EmoNeXt-Large [24] 0.0149 0.0065 0.0163 0.0207 0.0236 0.0228
ResidualMaskingNet [65] 0.0298 0.0307 0.0099 0.0251 0.0137 0.0280
Segm-VGG19 [98] 0.0186 0.0010 0.0174 0.0238 0.0221 0.0206
RAFDB DAN [102] 0.0319 0.0262 0.0017 0.0205 0.0246 0.0314
DDAMFN++ [110] 0.0013 0.0013 0.0002 0.0134 0.0117 0.0181
PosterV2 [50] 0.0326 0.0228 0.0054 0.0313 0.0166 0.0373
Next, we go one step further and quantitatively measure the impact of facial symmetry using our proposed score (Section 3.3). Table 2 summarizes these results averaged over all individuals $\mathfrak{I}$. Remember that a positive score corresponds to increased activations for increased symmetry. Further, we report the ratio of significant systematic changes over the set of individuals $\mathfrak{I}$ in Table 1 of the supplementary material. In most cases, we observe a significant impact of facial symmetry for all classifiers and expressions. Hence, this enables us to interpret the patterns we observe in Table 2 concerning the expressions and training datasets. We note the highest impact of $0.0373$ for surprise of PosterV2 [50]. This score indicates that, on average, over the complete surprise onset, increasing the symmetry by one step in our simulation increased the softmax output of PosterV2 by $3.7$ percentage points. However, while all 17 classifiers are significantly impacted by changes in facial symmetry, the effect size can still be small (see, for example, fear in Figure 5(c)).
We start with broad insights about the results in Table 2, before focusing on specific models: First, all $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)})$ contained in Table 2 are positive. Hence, expression classifiers show, on average, lower logit activations for decreased symmetry. This result provides interventional evidence for our previously stated hypothesis. Second, similar to Fig. 5, we see low scores for fear expressions irrespective of the classifier and dataset. However, this is likely due to the often lower activations for fear (Fig. 5 and supplementary material). The FLAME expression space may limit accurately modeling fear. The shift in model outputs is, nevertheless, significant. In contrast to fear, the overall high scores for happy, surprise, and angry suggest stronger changes in model behavior for these expressions.
Seen in Table 2, models trained on the same dataset often show similar $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)})$, likely due to the latent training data distribution [91]. For FER2013 [20] classifiers, facial symmetry has a lower impact on the disgust expression, excluding ResidualMaskingNet [65]. Similarly, models trained on this dataset display lower scores for sad and angry. We analyze different EmoNeXt [24] model sizes. Large being impacted the most, indicating higher capacity could consolidate biases in the training data. In contrast, DDAMFN++[12] shows a small effect size irrespective of the dataset. Visualizing the graphs at $t=1.0$ for different $s$ and the classification accuracies on the real-world data (both in the supplementary material), we assume that the model is likely overfitting the training data. Thus, low output activations result in smaller $\nabla_{s}$.
Nevertheless, we conclude that all analyzed classifiers show significant increases in output activations for higher facial symmetry, confirming our hypothesis. Given that we use interventions beyond the associational level, we verify a causal link between facial symmetry and expression classifiers’ behavior.
5 Limitations and Social Impact
Our work relies on the statistical shape model, which induces a domain shift. While many classifiers advertise the out-of-domain usage, e.g., [84, 83, 82, 65], and we additionally optimize the synthetic faces regarding model and expression (see Section 3.2). Nevertheless, other factors like camera angle could jointly influence the behavior. Secondly, while we focus on facial symmetry, other facial features, e.g., age or skin color, can impact classifier performance and should be considered in the future. Our framework cannot model the entire range of facial nerve damage effects on facial mimicry, as we do not account for all possible forms of facial asymmetry, e.g., synkinetic effects.
Regarding societal impact, our work investigates only existing expression classifiers. We move from the associational level to causal interventions to better understand how these black-box models operate. Such an approach could benefit other disciplines, primarily psychological and medical applications. We provide our experiments’ generative framework and evaluation code so that other researchers can simply evaluate their models.
6 Conclusion
Emotional expressiveness is crucial for communicating our internal state and for understanding other people [95, 76]. In this work, we investigate the impact of facial symmetry on 17 different expression classifiers trained on four different datasets [20, 47, 46, 53]. Extending empirical analysis, we try to answer an interventional question [4] by following insights from causal inference and explainability [73] and using an SCM (Fig. 1) together with a generative framework. We control expression and facial symmetry using a modified statistical shape model [48] to measure systematic changes with a proposed interpretable score.
We tested our hypothesis on real-world data of facial palsy patients and healthy probands using associational methods [73, 61, 9]. Here, we saw that facial palsy and the similarity between the face halves led to 33 out of 34 tests being significant. To verify these results in a more controlled manner, moving up the causal hierarchy, we employed our interventional framework to test the impact of symmetry on the models’ behavior. We observed that many classifiers, on average, decrease logit activations for lower facial symmetry.
While, in retrospect, our results align with the pre-specified intuition, we stress that our framework provides a structured way to test such hypotheses. Further, it could be extended to other features, e.g., age or skin color, given the controllable nature of statistical shape models. These insights could also be used to grade facial palsy or to correct the classier output for patients posthoc. Hence, we hope that our work can help researchers understand the prediction behavior of their trained expression classifiers beyond simple performance metrics.
6.0.1 Ethics Approval
Written informed consent was obtained from all participants, including displayed individuals. The ethics committee of the Jena University Hospital approved the study (No. 2019-1539).
6.0.2 Acknowledgement
Partially supported by Deutsche Forschungsgemeinschaft (DFG - German Research Foundation) project 427899908 BRIDGING THE GAP: MIMICS AND MUSCLES (DE 735/15-1 and GU 463/12-1).
6.0.3 Disclosure of Interests.
The authors have no competing interests to declare.
Appendix 0.A Structural Causal Models
Here we include a technical definition of structural causal models used in our work.
Definition 1 (Structural Causal Model in [60, Sec. 7.1.1] and [4, Def. 1])
A structural causal model (SCM) is defined as a 4-tuple $M=(U,V,F,P)$, where $U$ is a set of exogenous variables describing outside factors, $V=\{V_{1},...,V_{n}\}$ is the set of endogenous variables we measure in our model, $F=\{f_{1},...,f_{n}\}$ is a set containing functions $f_{i}$ that describe the functional relationships, and $P$ is a joint probability distribution over $U$. Further, each $V_{i}$ has a set of parents $PA_{i}$ that functionally determine $V_{i}$ together with some exogenous variables $U_{i}\subseteq U$. These parents $PA_{i}$ are a subset of $V\setminus\{V_{i}\}$. For settings $pa_{i}$ of parents $PA_{i}$ and $u_{i}$ of the exogenous variables $U_{i}$, $f_{i}$ determines the value $v_{i}=f_{i}(pa_{i},u_{i})$ of $V_{i}$.
Each causal model $M$ can be visualized as directed graphs. Here, each variable $V_{i}$ in $V$ defines a node, and we draw directed links from all parents $PA_{i}$ into $V_{i}$. Using such a model $M$, we can investigate questions of the following nature: given observed evidence, e.g., $V_{j}=v_{j}$, what is the probability of a statement $A$ happening? Further, performing a do-action on $V_{i}\in V$ is equivalent to removing the dependency $f_{i}$ and instead forcing $V_{i}$ to a constant value $x$. In other words, we set $F$ to $F_{x}$ with $F_{x}=\{f_{j}:V_{j}\neq V_{i}\}\cup\{V_{i}\leftarrow x\}$ [4].
Appendix 0.B Measuring Systematic Change - Significance Test
In Algorithm 1, we provide detailed pseudo code for our proposed shuffle hypothesis test regarding the significance of $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}})$. Further in Fig. 6, we visualize the estimated null-distribution as well as the originally measured score. We see that randomly shuffling observations along the symmetry axis results in a symmetrical distribution centered around zero, i.e., no systematic dependence on the facial symmetry. The original score, for the example in Fig. 6, is not typical for the estimated null-distribution leading to a low p-value. Table 3 contains the number of individuals per classifier and expression for which Algorithm 1 together with the Holm-Bonferroni correction [36] is significant. For this analysis, we perform the shuffle test with 10K iterations. All 17 classifiers show significant behavior changes concerning facial symmetry for all expressions and a majority of individuals.
Algorithm 1 Testing for statistical significance of $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}})$.
grid of predictions $\mathbb{F}_{\theta}^{(e)}(I_{\varphi^{(e)}}(s,t))$triangleright Gridsize is $S\times T$
integer K > 0 $\triangleright$ Number of Permutations
$\delta\in(0,1)$triangleright Significance Level
$p\leftarrow 0.0$
$\sigma_{orig.}\leftarrow\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}})$triangleright Estimate the original statistic
for $i\in\{1,...,K\}$ do
     $\mathbb{F}_{\theta}^{(e)}(I_{\varphi^{(e)}}^{(perm.)}(s,t))\leftarrow\texttt{permute}(\mathbb{F}_{\theta}^{(e)}(I_{\varphi^{(e)}}(s,t)),\texttt{axis}=0)$
$\triangleright$ Shuffle along Symmetry Axis
     $\sigma_{perm.}\leftarrow\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}}^{(perm.)})$
     if $|\sigma_{perm.}|>|\sigma_{orig.}$then $\triangleright$ Absolutes because our statistic is two sided
         $p\leftarrow p+\nicefrac{{1}}{{K}}$triangleright Increment the $p$-value
     end if
end for
if $p<\delta$ then
     return $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}})$ is significant.
else
     return $\hat{\mathcal{S}}(\mathbb{F}_{\theta}^{(e)}|I_{\varphi^{(e)}})$ is not significant.
end if
Figure 6: Using the shuffle test, outlined in Algorithm 1, we plot the resulting scores for $100000$ permutations in a histogram. Our tests use a significance threshold of $p<(0.05)$. The original model score for a single individual, shown as a red dashed line, lies clearly outside the computed null distribution and is thus significant.
Table 3: We report how many of the 200 individuals, using the Holm-Bonferroni [36] corrected p-values, have been significant ($p<0.05$). We can see that the majority of all results are significant, confirming our hypothesis that facial symmetry impacts the internal decision rules.
Dataset Model Angry Disgust Fear Happy Sad Surprise
AffectNet7 DAN [102] 200 200 200 200 200 200
DDAMFN++ [110] 200 200 200 200 200 200
HSEmotion [82] 200 200 200 200 200 200
PosterV2 [50] 200 199 200 200 200 200
AffectNet8 DAN [102] 200 200 200 200 200 200
DDAMFN++ [110] 200 200 198 200 200 200
HSEmotion [82] 200 200 200 200 200 200
PosterV2 [50] 200 200 200 200 200 200
FER2013 EmoNeXt-Tiny [24] 200 200 200 200 200 200
EmoNeXt-Small [24] 199 194 200 200 200 200
EmoNeXt-Base [24] 163 137 200 200 200 200
EmoNeXt-Large [24] 133 187 196 195 198 187
ResidualMaskingNet [65] 200 194 199 200 200 200
Segmentation-VGG19 [98] 199 148 200 148 200 129
RAFDB DAN [102] 200 199 200 200 200 200
DDAMFN++ [110] 200 200 200 200 200 200
PosterV2 [50] 200 194 200 200 195 200
Appendix 0.C Additional Details Experiment 1
This section gives an overview of the prediction accuracy of all 17 expression classifiers achieved on our real-world data, consisting of healthy probands and patients with unilateral facial palsy. Further, we detail the hyperparameter choices in our experiments regarding the associational methods to infer whether a causal link exists between facial symmetry and model prediction behavior. Lastly, we include some additional visualizations regarding the symmetry features.
0.C.1 Real-World Prediction Accuracy
Table 4: We evaluated each classifier on the faces of the healthy probands and patients with unilateral facial palsy mimicking the happy facial expression. Low accuracy is displayed in a darks shade, and high accuracy is displayed in a light shade. Models trained on FER2013 especially seem to work well on our data set. Models trained on RAFDB seem to be less suitable. Further, we provide the mean accuracy of all models per data set (with at least one correct classification).
Dataset Model $s=0.0$ $s=0.5$ $s=1.0$ Probands Patients
AffectNet7 DAN [102] 57.50% 99.00% 100.00% 94.44% 62.82%
DDAMFN++ [110] 0.00% 0.00% 19.50% 0.35% 0.20%
HSEmotion [82] 96.00% 100.00% 100.00% 0.00% 0.00%
PosterV2 [50] 87.00% 100.00% 100.00% 97.92% 61.23%
Average 80.17% 99.67% 79.88% 64.24% 41.42%
AffectNet8 DAN [102] 0.00% 30.50% 88.50% 91.32% 45.92%
DDAMFN++ [110] 0.00% 8.00% 45.00% 0.35% 6.56%
HSEmotion [82] 0.00% 0.00% 97.50% 0.00% 0.00%
PosterV2 [50] 0.00% 20.50% 99.00% 89.93% 36.58%
Average 0.00% 19.67% 82.50% 60.53% 29.69%
FER2013 EmoNeXt-Base [24] 10.50% 71.00% 97.50% 99.65% 70.38%
EmoNeXt-large [24] 28.50% 77.50% 100.00% 98.26% 63.42%
EmoNeXt-small [24] 10.50% 68.50% 100.00% 98.96% 66.40%
EmoNeXt-tiny [24] 0.00% 12.50% 87.00% 97.92% 58.45%
ResidualMaskingNet [65] 33.50% 89.00% 100.00% 92.36% 59.05%
Segmentation-VGG19 [98] 2.00% 39.50% 99.00% 97.22% 31.81%
Average 17.00% 59.67% 97.25% 97.40% 58.25%
RAFDB DAN [102] 30.00% 65.50% 94.50% 32.29% 54.27%
DDAMFN++ [110] 69.00% 99.00% 100.00% 90.97% 76.34%
PosterV2 [50] 25.50% 80.00% 100.00% 8.68% 38.17%
Average 41.50% 81.50% 98.17% 43.98% 56.26%
Total Average 40.91% 64.03% 89.85% 72.71% 48.77%
We are interested in the overall prediction accuracy of the model on our real-world data set consisting of $36$ healthy probands and $36$ patients with unilateral facial palsy. Both were instructed to mimic a happy expression. The probands repeated the information four times in two sessions, yielding 288 images. The patients followed the same instruction video during a ten-day biofeedback training at the hospital. They also repeated the exercise four times during a session on the first, third, and last day of therapy. An additional fourth session was offered after six months but was not followed up by some patients. Thus, we obtained 503 images for the patients.
In Table 4, we display the prediction accuracy of the happy emotion. We see strong differences per model and group. Therefore, we also denote the average accuracy per model and dataset to understand how we can see a particular trend per dataset. As expected and shown in the main paper, the performance of the models degrades for images that contain some form of facial asymmetry (either simulated at $s=0.0$, $s=0.5$, or actual facial palsy). Thus, we assume that facial symmetry is the underlying cause impacting the internal decision rules of the black box classifiers. We also see that the DDAMFN++ model trained on the AffectNet similarly performs worse on our real-world data than on the synthetic data we use for our intervention framework. Interestingly, the RAFDB-provided checkpoints seem more robust, at least in the case of happy.
Given that we also follow a similar experimental setup as in FER2013, models trained on it have the best performance on our data, observable in the table. Several reasons could be involved; either mimicry and natural facial expression have some inherent differences, the model source on public data (and human annotated) cannot differentiate, or the impact of confounding factors like camera pose. Lighting ensures that the models focus more on facial expressions.
Models such as PosterV2 perform well in our synthetic framework (likely due to the optimized expression parameters). Still, they seemed to overfit on the training data RAFDB as they performed worse on the probands but somehow better on the patients.
0.C.2 Feature Attribution Hyperparameter Choices
Our main experiment 2 tests the statistical dependence between expression classifier outputs and facial symmetry. We focus on the happy logit and find that most models change their behavior significantly for variations in facial symmetry. We employ the feature attribution method described in [73] toward this goal. This method frames supervised learning as an SCM [60] and tests whether network predictions and a pre-defined feature (facial symmetry) are conditionally independent given the reference annotation. If we have to discard this null hypothesis, we know that the classifier output values vary significantly for changes in the investigated feature. This procedure is motivated by Reichenbach’s common cause principle [70].
Clearly, the choice of conditional independence test is an important hyperparameter choice to ensure that the results are reliable. Further, Shah and Peters [86] prove that there is no optimal test that can control type-I errors, i.e., false positives, irrespective of the joint latent distribution in the non-parametric case. Because we have no knowledge about the joint distribution of all variables important in our analysis, we are exactly in the non-parametric case. Here, we follow previous work [72, 62, 61, 9] and select multiple non-linear tests. Specifically, we select conditional HSIC [27], CMIknn [79], and FCIT [11]. We consider the result from all three tests and report the majority decision [72].
The selected conditional independence tests themselves have different hyperparameter choices. First, for conditional HSIC [27], we have to select a suitable kernel function. We follow the suggestion of the authors and select the common radial basis functions kernel. Additionally, we use the heuristic by Gretton et al. [31] to approximate suitable kernel widths for all of our three variables. Second, similarly for CMIknn [79], we follow the suggested hyperparameter settings. Specifically, we set $k_{perm.}$, i.e., the neighborhood size, to five and use ten percent of the data to estimate the conditional mutual information ($k_{CMI}=0.1\cot n$ for n data points). Lastly, for FCIT [11], we again follow the suggestions by the authors. In other words, we set the number of data permutations to eight and use ten percent of the data to calculate the test statistic.
0.C.3 Additional Visualizations Regarding Logit Activations
Following previous work [9], we visualize the difference in the happy logit behavior between the healthy probands and facial palsy patients. Fig. 7 contains these results split between the training datasets of the 17 models we investigate in this work. However, these are associational investigations, i.e., of the first level of the PCH [4]. In other words, we do not isolate changes in facial symmetry from confounding factors and it is highly likely that other features correlate with the presence of facial palsy. Hence, while we observe changes in classifier behavior on real data, our interventional investigation is more reliable and provides actionable insights.
Nevertheless, Fig. 7 shows a decrease in happy activations for most models. This is congruent with the aggregated performance results in Table 4. Further, these results are in line with our insights gained using our interventional framework: asymmetry results in lower activations for the happy class. Interestingly, we observe a slight deviation for models trained on the RAFDB. Here DAN [102], and PosterV2 [50] show higher activations and improved performance. Nonetheless, both models still struggle with facial palsy patients and are outperformed by DDAMMFN++ [110] trained on the same dataset.
Additionally, we also visualize the results for the continuous LPIPS [109] symmetry. For regressing the mean and standard deviation, we use a window regression approach as described in [61]. We display these visualizations in Fig. 8.
Overall, we observe for most models a decrease in logit activations for decreasing facial symmetry. Hence, these results are congruent with the findings made in the main paper and Fig. 7. Furthermore, we again observe a very small effect size for DDAMFS++ [110] for both features. These findings are in agreement with the noted performance in Table 4.
Nevertheless, we want to highlight two additional observations: First, in Figure 8(d), we observe an unexpected increase in activations. While these are associational insights, i.e., there are many possible reasons, these increases are also visible in Figure 7(d). Second, while for most models in Fig. 8, we observe a decrease in logit activations for lower facial symmetry, we note a smaller increase again for the most asymmetric faces.
(a) Shift in output behavior for classifiers trained on AffectNet7 [53] with respect to facial palsy.
(b) Shift in output behavior for classifiers trained on AffectNet8 [53] with respect to facial palsy.
(c) Shift in output behavior for classifiers trained on FER2013 [20] with respect to facial palsy.
(d) Shift in output behavior for classifiers trained on RAFDB [47, 46] with respect to facial palsy. Note that we find the behavior shift for the DAN [102] model is not significant.
Figure 7: We follow [9] and visualize the differences in the classifiers happy logit distribution for healthy probands and facial palsy patients. Here 7(a) - 7(d) contain models trained on the indicated dataset respectively.
(a) Shift in output for classifiers trained on AffectNet7 [53] with respect to LPIPS [109] symmetry.
(b) Shift in output for classifiers trained on AffectNet8 [53] with respect to LPIPS [109] symmetry.
(c) Shift in output for classifiers trained on FER2013 [20] with respect to LPIPS [109] symmetry.
(d) Shift in output for classifiers trained on RAFDB [47, 46] with respect to LPIPS [109] symmetry.
Figure 8: We follow [61, 9] and regress the shift in the classifiers happy logit distribution for measured LPIPS [109] symmetry scores of healthy probands and facial palsy patients. Here 8(a) - 8(d) contain models trained on the indicated dataset respectively. Note that higher LPIPS corresponds to lower symmetry [109].
Appendix 0.D Additional Details Experiment 2
This section details the behavior analysis of the 17 expression classifiers on our synthetic intervention data. We start with setup and information about the facial expression optimization before displaying the sampled individuals. Afterward, we display the facial expressions achieved during the optimization per classifier for an individual. Finally, we visualize the resulting activation surfaces.
Table 5: Using our intervention framework, we optimized each expression classifier $\mathbb{F}_{\theta}$ using logit activation for each of the six base emotions. We display the average logit activation per model and emotion. Low activation is displayed in a darks shade, and high activation is displayed in a light shade. We observe that fear has a generally low activation, indicating that the models have issues classifying fear or that the FLAME expression cannot model fear.
Dataset Model Angry Disgust Fear Happy Sad Surprise
AffectNet7 DAN [102] 0.862 0.853 0.446 0.842 0.702 0.917
DDAMFN++ [110] 0.356 0.331 0.136 0.220 0.233 0.292
HSEmotion [82] 0.915 0.913 0.403 0.979 0.824 0.954
PosterV2 [50] 0.835 0.931 0.505 0.950 0.747 0.931
AffectNet8 DAN [102] 0.776 0.805 0.416 0.464 0.732 0.881
DDAMFN++ [110] 0.237 0.211 0.122 0.228 0.212 0.316
HSEmotion [82] 0.595 0.759 0.349 0.340 0.590 0.826
PosterV2 [50] 0.814 0.911 0.499 0.666 0.725 0.928
FER2013 EmoNeXt-Tiny [24] 0.400 0.083 0.269 0.508 0.278 0.821
EmoNeXt-Small [24] 0.727 0.088 0.342 0.752 0.345 0.885
EmoNeXt-Base [24] 0.548 0.076 0.432 0.720 0.465 0.884
EmoNeXt-Large [24] 0.902 0.352 0.644 0.886 0.829 0.862
ResidualMaskingNet [65] 0.959 0.995 0.500 0.884 0.581 0.997
Segmentation-VGG19 [98] 0.818 0.067 0.725 0.976 0.846 0.919
RAFDB DAN [102] 0.991 0.877 0.088 0.874 0.947 0.999
DDAMFN++ [110] 0.077 0.070 0.013 0.585 0.808 0.771
PosterV2 [50] 0.993 0.996 0.312 0.982 0.987 1.000
0.D.1 Classifier Facial Expression Optimization
Our experiments optimized each classifier $\mathbb{F}_{\theta}$ regarding the six base emotions. Therefore, we report the average logit activation per model and emotion reached in Table 5. We can observe several interesting properties in the logit activation. First, not all models can reach high logit activation based on facial expression changes. This indicates that models also leverage other facial information while classifying facial expressions. Furthermore, we observe that fear has a low activation among all classifiers except SegmentationVgg19 [98]. The surprise facial expression has a high activation among all classifiers, whereas DDAMFN++[110] is the sole outlier; overall, reached activation is low.
0.D.2 Individuals
We provide an overview of all created individuals in Fig. 9. The data can be downloaded here:https://doi.org/10.6084/m9.figshare.27074587.v1. All resemblance to existing people is not intended and could only result from the underlying FLAME geometry model [48] and the texture from the BaselFaceModel [59].
Figure 9: 200 individual population $\mathfrak{I}$
0.D.3 Average Facial Expression
Together with the reached logit activations, see Table 5, we are interested in the resulting facial expression. These should depict the internal representation of the respective emotion and give insight into what each classifier assumes. Furthermore, we assume the underlying base dataset influences the expression.
0.D.3.1 Average Facial Expression - Per Dataset
Using our generative facial expression network, we can now create a representation of how different classifiers represent the underlying training dataset. This means the expression vectors of all 200 individuals per dataset and model are averaged and shown in Fig. 10. This visualization gives an intuitive feeling about the underlying facial expression per FER benchmark [53, 47, 46, 20]. Looking at the expression columns, we see that all interpretations of a face are slightly different. For example, for angry, the mouth frowning angles are different. For disgust the mouth is slightly opened compared to angry. For fear we can clearly see that raising the eye brows is common. The happy expression varies either with a wider grin or the opening state of the eye. The sad expression varies in the intensity of the frowning, but the eyebrows are not activated by the corrugator muscle. Also, the eyes are generally closed. For the surprise expression, we can see wide-open eyes and raised eyebrows in the shared interpretation.
Even though they are similar in their visual state, the intensity and expressiveness are different per model and could be the underlying reason for differences in the model architectures or the data used in the benchmark.
Figure 10: Average Facial Expression used for classification based on the underlying training dataset.
0.D.3.2 Average Facial Expression - AffectNet7
Fig. 11 contains the average facial expressions for models trained on AffectNet7 [53].
Figure 11: The average facial expressions for models trained on AffectNet7 [53]
0.D.3.3 Average Facial Expression - AffectNet8
Fig. 12 contains the average facial expressions for models trained on AffectNet8 [53].
Figure 12: The average facial expressions for models trained on AffectNet8 [53]
0.D.3.4 Average Facial Expression - FER2013
Fig. 13 contains the average facial expressions for models trained on FER2013 [20].
Figure 13: The average facial expressions for models trained on FER2013 [20]
0.D.3.5 Average Facial Expression - RAFDB
Fig. 14 contains the average facial expressions for models trained on RAFDB [47, 46].
Figure 14: The average facial expressions for models trained on RAFDB [47, 46]
0.D.4 Model Activation Surfaces
The main paper shows that we use a finite grid over $\mathfrak{T}$ and $\mathfrak{S}$ to compute the FAIS score. Given that we only highlighted the final time step $t=1.0$, we show here the full logit activation surfaces used to compute our score.
(a) HSEmotion [82]
(b) DAN [102]
(c) DDAMFN++ [110]
(d) PosterV2 [50]
Figure 15: AffectNet7 [53]
(a) HSEmotion [82]
(b) DAN [102]
(c) DDAMFN++ [110]
(d) PosterV2 [50]
Figure 16: AffectNet8 [53]
(a) ResidualMaskingNet [65]
(b) SegmentationVGG19 [98]
(c) EmoNeXt-Tiny [24]
(d) EmoNeXt-Small [24]
(e) EmoNeXt-Base [24]
(f) EmoNeXt-Large [24]
Figure 17: FER2013 [20]
(a) DAN [102]
(b) DDAMFN++ [110]
(c) PosterV2 [50]
Figure 18: RAFDB [47, 46]
0.D.5 Model Symmetry Impact
The main paper shows that we compute the interpretable asymmetry score using a finite grid over $\mathfrak{T}$ and $\mathfrak{S}$. Given that we only highlighted the final time step $t=1.0$, we show here the full logit activation surfaces used to compute our score.
(a) HSEmotion [82]
(b) DAN [102]
(c) DDAMFN++ [110]
(d) PosterV2 [50]
Figure 19: AffectNet7 [53]
(a) HSEmotion [82]
(b) DAN [102]
(c) DDAMFN++ [110]
(d) PosterV2 [50]
Figure 20: AffectNet8 [53]
(a) ResidualMaskingNet [65]
(b) SegmentationVGG19 [98]
(c) EmoNeXt-Tiny [24]
(d) EmoNeXt-Small [24]
(e) EmoNeXt-Base [24]
(f) EmoNeXt-Large [24]
Figure 21: FER2013 [20]
(a) DAN [102]
(b) DDAMFN++ [110]
(c) PosterV2 [50]
Figure 22: RAFDB [47, 46]
0.D.6 Local Explanations - Saliency Maps
We aim to understand the impact of facial asymmetry globally; local explanations via saliency maps still offer insights but require human interpretation. We use an occlusion-based interpretation approach [107] for the ground truth and predicted label using the average emotion simulated with the default identity.
0.D.6.1 Local Explanations - AffectNet7
The saliency maps indicate that independent of the predicted or the ground truth label, the majority impact is only on one side of the face. This supports our global observation that facial symmetry has a strong impact on the model behavior.
(a) Model focus based on the ground truth label.
(b) Model focus based on the predicted truth label.
Figure 23: The occlusion-based saliency maps for models trained on AffectNet7 [53]
0.D.6.2 Local Explanations - AffectNet8
The saliency maps indicate that independent of the predicted or the ground truth label, the majority impact is only on one side of the face. This supports our global observation that facial symmetry has a strong impact on the model behavior.
(a) Model focus based on the ground truth label.
(b) Model focus based on the predicted truth label.
Figure 24: TThe occlusion-based saliency maps for models trained on AffectNet8 [53]
0.D.6.3 Local Explanations - FER2013
The saliency maps indicate that independent of the predicted or the ground truth label, the majority impact is only on one side of the face. This supports our global observation that facial symmetry has a strong impact on the model behavior.
(a) Model focus based on the ground truth label.
(b) Model focus based on the predicted truth label.
Figure 25: The occlusion-based saliency maps for models trained on FER2013 [20]
0.D.6.4 Local Explanations - RAFDB
The saliency maps indicate that independent of the predicted or the ground truth label, the majority impact is only on one side of the face. This supports our global observation that facial symmetry has a strong impact on the model behavior.
(a) Model focus based on the ground truth label.
(b) Model focus based on the predicted truth label.
Figure 26: The occlusion-based saliency maps for models trained on RAFDB [47, 46]

### v2

@@
 The state-of-the-art for automatically classifying the six base emotions [23] is achieved by end-to-end trained black box neural networks [65, 82, 110, 102, 24, 17].
-However, it remains unclear how the internal decision-making processes of these models respond to out-of-distribution inputs that were not accounted for during training.
+However, it remains unclear how the internal decision-making processes of these models respond to out-of-distribution inputs due to likely unbalanced training data.
 Specifically, we observe a performance degradation when classifying facial expressions in individuals with unilateral facial palsy, a condition impairing the ability to produce symmetrical facial expressions due to underlying nerve damage.
@@
 Causal inference tries to answer causal questions from data [64].
-This includes interventions, i.e., additional experiments, and purely observational data.
+This includes interventions, i.e., additional experiments and purely observational data.
 Importantly, causal questions can be categorized into a hierarchy.
-This so-called Pearl’s causal hierarchy (PCH) [4] consists of the following three levels ordered by increasing difficulty: associational, interventional, and counterfactual questions.
-The latter two are analyzed using the do-operator [60], which changes a variable to a constant value, e.g., we would write $do(\text{Facial Symmetry}:=s)$ for the variable facial symmetry.
+This so-called Pearl’s causal hierarchy (PCH) [4] consists of the three levels ordered by increasing difficulty: associational, interventional, and counterfactual questions.
+The latter two are analyzed using the do-operator [60], which changes a variable to a constant value, e.g., we write $do(\text{Facial Symmetry}:=s)$ for the variable facial symmetry.
 Furthermore, framing data-generating processes and complex interactions of our physical reality as directed graphs enables us to precisely define and investigate the underlying causal mechanisms [60, 64].
@@
 Our study is limited to the six base emotions [23], omitting neutral and contempt for comparability.
-Compared to other research that requires training [43, 15], we can asses 17 existing expression classifiers.
+We can assess existing expression classifiers, whereas other research requires training [43, 15].
 Please note that our selection of models is not intended to be comprehensive or representative of all possible classifiers but as a first case study.
-We focus on a subset of models that provide code and model weights, which are likely to be applied more frequently in out-of-domain scenarios like medicine and psychology.
+We focus on a subset of models that provide code and model weights, which are likely to be applied in out-of-domain scenarios like medicine and psychology.
 To test our hypothesis, we perform two groups of experiments.
 First, we investigate if a symmetry bias exists for real-world data inside classifiers using associational methods [73, 71, 61, 7].
-Second, moving up on the causal hierarchy [4], we connect facial symmetry and model output using our synthetic intervention framework and score.
+Second, moving up on the causal hierarchy [4], we link facial symmetry and model output using our synthetic intervention framework.
 4.1 Experiment 1: Observations on Real-World Data
@@
 4.
-We measure an average classification accuracy of $97.40\%$ for probands and $58.25\%$ for patients among the FER2013 models (due to the similar recording setup), indicating the decreased performance under facial palsy.
-A more fine-grained analysis can be found in the supplementary material.
+We measure an average classification accuracy of $97.40\%$ for probands and $58.25\%$ for patients among the FER2013 models, indicating the decreased performance under facial palsy.
+A more detailed analysis can be found in the supplementary material.
 DAN [102] DDAMFN++ [110] HSEmotion [82] PosterV2 [50] DAN [102] DDAMFN++ [110] HSEmotion [82] PosterV2 [50] EmoNeXt-Base [24] EmoNeXt-Large [24] EmoNeXt-Small [24] EmoNeXt-Tiny [24] ResidualMaskingNet [65] Segmentation-VGG19 [98] DAN [102] DDAMFN++ [110] PosterV2 [50]
@@
 Table 1: Significance results ($p<0.01\rightarrow$ ) of [73] on our data for three symmetry features.
-We investigate happy logits with respect to the binary facial palsy state and LPIPS [108] similarity between the face sides for real-world images of facial palsy patients and healthy probands.
+We analyze happy logits regarding the binary facial palsy state and LPIPS [108] similarity between the face sides for facial palsy patients and healthy probands images.
 Following related work [9, 73, 61], we denote all significant behavior changes in Table 1 using the majority decision of three conditional independence tests [27, 79, 11] (for detailed hyperparameters see supplementary material).
@@
 However, the continuous symmetry measure LPIPS [108] indicates the same behavior changes.
-These results provide first evidence for our hypothesis that expression classifiers are biased toward facial symmetry, especially concerning downstream applications.
-Furthermore, we provide more qualitative investigations of the output changes in the supplementary material.
+These results provide evidence for our hypothesis that expression classifiers are biased toward facial symmetry, especially concerning downstream applications.
+We provide more qualitative investigations of the output changes in the supplementary material.
 For most classifiers, we observe, on average, a decrease in activations.
-Given the performance decrease, this observation is expected and indicates uncertainty in predicting the happy class.
-However, these are associational investigations [4], i.e., we do not isolate changes in facial symmetry from confounding factors.
-Hence, while we observe changes in classifier behavior on real-world data, our following interventional investigation in Section 4.2 provides more reliable, actionable insights.
+Given this decrease, this observation is expected and indicates uncertainty in predicting the happy class.
+However, these are associational investigations [4], i.e., we cannot isolate changes due to only facial symmetry.
+Hence, while we observe changes in classifier behavior on real-world data, our interventional investigation in Section 4.2 provides more reliable, actionable insights.
 4.2 Experiment 2: Synthetic Facial Symmetry Interventions
@@
 5.
-First, we observe that facial symmetry impacts each expression’s logit activation for the models irrespective of the dataset.
+We observe that facial symmetry impacts each expression’s logit activation for the models irrespective of the dataset.
 The ResidualMaskinNet [65] trained on FER2013 [20] does not reach high activations for fear and sad.
@@
 PosterV2 [50] 0.0326 0.0228 0.0054 0.0313 0.0166 0.0373
-Next, we go one step further and quantitatively measure the impact of facial symmetry using our proposed score (Section 3.3).
+We go one step further and quantitatively measure the impact of facial symmetry using our proposed score (Section 3.3).
 Table 2 summarizes these results averaged over all individuals $\mathfrak{I}$.
-Remember that a positive score corresponds to increased activations for increased symmetry.
+A positive score corresponds to increased activations for increased symmetry.
 Further, we report the ratio of significant systematic changes over the set of individuals $\mathfrak{I}$ in Table 1 of the supplementary material.
 In most cases, we observe a significant impact of facial symmetry for all classifiers and expressions.
-Hence, this enables us to interpret the patterns we observe in Table 2 concerning the expressions and training datasets.
+This enables us to interpret the patterns we observe in Table 2 concerning the expressions and training datasets.
 We note the highest impact of $0.0373$ for surprise of PosterV2 [50].
@@
 5 Limitations and Social Impact
-Our work relies on the statistical shape model, which induces a domain shift.
-While many classifiers advertise the out-of-domain usage, e.g., [84, 83, 82, 65], and we additionally optimize the synthetic faces regarding model and expression (see Section 3.2).
-Nevertheless, other factors like camera angle could jointly influence the behavior.
-Secondly, while we focus on facial symmetry, other facial features, e.g., age or skin color, can impact classifier performance and should be considered in the future.
-Our framework cannot model the entire range of facial nerve damage effects on facial mimicry, as we do not account for all possible forms of facial asymmetry, e.g., synkinetic effects.
-Regarding societal impact, our work investigates only existing expression classifiers.
+Our work relies on the statistical shape model, inducing a domain shift and limiting possible expressions.
+Some classifiers advertise the out-of-domain usage, e.g., [84, 83, 82, 65], and we optimize the synthetic faces regarding model and expression (see Section 3.2).
+Factors like camera angle could jointly influence the behavior.
+Secondly, other facial features, e.g., age or skin color, could impact classifier performance and should be considered.
+In our current framework, we cannot account for all possible forms of facial asymmetry, e.g., synkinetic effects.
+Regarding societal impact, we investigate existing expression classifiers only.
 We move from the associational level to causal interventions to better understand how these black-box models operate.
-Such an approach could benefit other disciplines, primarily psychological and medical applications.
-We provide our experiments’ generative framework and evaluation code so that other researchers can simply evaluate their models.
+This could benefit other disciplines, primarily psychological and medical applications.
+We provide our experiments’ framework and evaluation code so that researchers can evaluate their models.
 6 Conclusion
@@
 We control expression and facial symmetry using a modified statistical shape model [48] to measure systematic changes with a proposed interpretable score.
-We tested our hypothesis on real-world data of facial palsy patients and healthy probands using associational methods [73, 61, 9].
+We tested our hypothesis on real-world data using associational methods [73, 61, 9].
 Here, we saw that facial palsy and the similarity between the face halves led to 33 out of 34 tests being significant.
-To verify these results in a more controlled manner, moving up the causal hierarchy, we employed our interventional framework to test the impact of symmetry on the models’ behavior.
+To verify these results in a controlled manner, moving up the causal hierarchy, we employed our interventional framework to test the impact of symmetry on the models’ behavior.
 We observed that many classifiers, on average, decrease logit activations for lower facial symmetry.
@@
 6.0.1 Ethics Approval
-Written informed consent was obtained from all participants, including displayed individuals. The ethics committee of the Jena University Hospital approved the study (No. 2019-1539).
+Written consent was obtained from all participants. The Jena University Hospital ethics committee approved the study (No. 2019-1539).
 6.0.2 Acknowledgement
 Partially supported by Deutsche Forschungsgemeinschaft (DFG - German Research Foundation) project 427899908 BRIDGING THE GAP: MIMICS AND MUSCLES (DE 735/15-1 and GU 463/12-1).
-6.0.3 Disclosure of Interests.
-The authors have no competing interests to declare.
 Appendix 0.A Structural Causal Models
@@
 Definition 1 (Structural Causal Model in [60, Sec. 7.1.1] and [4, Def. 1])
-A structural causal model (SCM) is defined as a 4-tuple $M=(U,V,F,P)$, where $U$ is a set of exogenous variables describing outside factors, $V=\{V_{1},...,V_{n}\}$ is the set of endogenous variables we measure in our model, $F=\{f_{1},...,f_{n}\}$ is a set containing functions $f_{i}$ that describe the functional relationships, and $P$ is a joint probability distribution over $U$.
+A structural causal model (SCM) is defined as a 4-tuple $M=(U,V,F,P)$, where $U$ is a set of exogenous variables describing outside factors, $V=\{V_{1},...,V_{n}\}$ is the set of endogenous variables we measure in our model, $F=\{f_{2},...,f_{n}\}$ is a set containing functions $f_{i}$ that describe the functional relationships, and $P$ is a joint probability distribution over $U$.
 Further, each $V_{i}$ has a set of parents $PA_{i}$ that functionally determine $V_{i}$ together with some exogenous variables $U_{i}\subseteq U$.